A typical gas thermometer is calibrated between temperatures 90C and + 105C and at pressures 0.750 atm and 1.60 atm respectively.


Calculate the pressure at the freezing point of water.


Grade:Upto college level

1 Answers

Navneet kumar
10 years ago
Let X be any thermometric property that varies linearly with temperature. X can be the length of a metal rod, the height of a column of mercury, the electrical resistance of a wire, the pressure of a gas, etc. X ? T Since X has linear variation with T, a plot of X versus T is a straight line: X = C · T + X0 X0 is the value of X at T=0. In the Celsius scale And C is the slope = (1.60- . 75 )/ (105 - 90) = .85 / 15 to find out X0 = X - CT = 1.60 - (.85/ 15 ) x 105 = A suction of 4.95 atm